# DeltaPhi: Learning Physical Trajectory Residual for PDE Solving

### v1

@@ 3.2.2 Auxiliary Trajectory Sampling
Inference stage. During inference, given the input function $a^{test}$, we select the most similar auxiliary trajectory from the training set:
$\mathcal{F}^{ret}(a^{test},\mathcal{T})=\texttt{argmax}_{j}\,\texttt{Sim}(a^{test},a_{j}).$ (7)
We empirically found the model performance is robust to the choice of auxiliary trajectories. In Section A.3, we experiment with randomly sampled auxiliary trajectories from the top ten similarity scores and find the standard deviation across five repeating evaluations is quite small. This indicates that the residual neural operator effectively learns the residual operator mapping for the same $a_{i}$ with varied auxiliary trajectories. We further present the trajectory sampling for varied resolution fields in Section C.3. Section C.4 provides the cost analysis of the trajectory sampling.
3.2.3 Auxiliary Trajectory Integration
As aforementioned, the residual operator $\mathcal{G}^{\Delta}$ requires additional auxiliary function $a_{k_{i}}$ as input. In specific implementation, $a_{i}$ and $a_{k_{i}}$ could be straightly concatenated along their channel dimension:
@@ 3.2.3 Auxiliary Trajectory Integration
where $\oplus$ represents the concatenation operation of two vectors. Then the concatenated vector $a_{input}$ is feeded into the neural operator network:
@@
where $P$ is the fully connected encoder in the operator network (Equation 3), and $v_{0}$ represents the first latent function in the hidden space.
Integration of auxiliary trajectory with variant resolution. In cases where the auxiliary input function $a_{k_{i}}$ has a different resolution from the input function $a_{i}$, we address this discrepancy via field interpolation techniques. We present the details in Section C.2.
Integration of auxiliary trajectory with additional related information. The residual operator mapping can be challenging to learn due to the variability of output residuals for different auxiliary samples. To mitigate such difficulty, we introduce additional information related to $a_{k_{i}}$ as inputs for residual neural operators:
  • For time-series PDE solving, e.g. Navier-Stokes, we only input the last few time steps of $a_{k_{i}}$.
  • We use the corresponding $u_{k_{i}}(x)$ of $a_{k_{i}}(x)$ as the input.
  • We utilize calculated relation metrics, e.g. similarity between $a_{i}$ and $a_{k_{i}}$, as inputs.
Our experimental results (Section A.2) indicate that this customized input strategy indeed improves the performance of residual neural operators.
3.2.4 Residual Operator Learning
To train the parameterized surrogate model $\mathcal{G}_{\theta}^{\Delta}$ of $\mathcal{G}^{\Delta}$ using the residual training set $\mathcal{T}^{res}$, we formulate the following optimization problem to determine the optimal parameter by minimizing the cost function:
$\ell=\sum_{i=1}^{N}\mathcal{L}(\mathcal{G}_{\theta}^{\Delta}(a_{i},a_{k_{i}}),u_{i}-u_{k_{i}}),$ (10)
where $\mathcal{L}$ is the cost function for assessing the discrepancy between function residuals.
The optimization of the physical trajectory residual depicted in Equation 10 can be effectively implemented using existing learning frameworks [2, 11]. This involves introducing a trajectory residual connection on the explicit output of the base operator network. Specifically, we add the auxiliary solution $u_{k_{i}}$ to the predicted residual from the final output of the operator network:
@@ 3.2.4 Residual Operator Learning
where $Q$ is the last fully connected projector of the operator network, $v_{l}$ is the last latent function in hidden space produced by $\sigma(W_{l}+\mathcal{K}_{l})$ shown in Equation 3.
During inference, given input function $a^{test}$, we select the auxiliary trajectory $(a_{k_{t}}(x),u_{k_{t}}(x))$ from the training set $\mathcal{T}$ using the sampling function $\mathcal{F}^{ret}$.
3.2.5 Discussion
Compared to direct operator learning, residual operator learning presents the following advantages.
Mitigation of distribution bias. Residual operator learning allows us to control the training label distribution, thereby mitigating the data distribution bias. We posit that the similarity between input functions and the corresponding output functions’ similarity are positively correlated. This correlation is validated empirically in non-chaotic systems, as demonstrated in Darcy Flow experiments (Figure 4). Consequently, the distribution of label residuals can be effectively controlled by adjusting the selection range ($K$ in Equation 6) of the sampling function $\mathcal{F}^{ret}$. This adjustment allows the construction of sufficiently diverse training labels, thereby alleviating distribution bias in restricted data scenarios and enhancing the generalization capabilities of neural operators.
Prevent overfitting. In scenarios with limited training data, neural networks tend to memorize the training set, leading to significant overfitting. Residual learning effectively addresses this issue by modifying the training labels through the use of various auxiliary samples in each training epoch, thus reducing the risk of overfitting.
Incorporation of explicit physical priors. The auxiliary sample acts as a non-parametric physical prior. As shown in previous work [12], involving external memory components benefits the model’s generalization capability. Moreover, the explicit external memory allows specific needs during inference, such as handling boundary conditions and high resolution samples.
@@ 4 Experiment
This section provides the experiments on various PDEs, resolution generalization problem, and statistical analysis of residual operator learning.  Appendix A presents additional ablation studies.
4.1 Residual Learning vs. Direct Learning on Various PDEs
This section validates residual operator learning on various PDEs. The experiment covers both irregular domain problems and regular domain problems. In addition, we test the performance improvement under variant training data amount. To ensure fair comparisons, all baselines and our counterpart models are compared with exactly same hyperparameters, trainable weight amount, optimizer configuration, and training step.
Irregular domain problems. We validate the effectiveness of the proposed approach on irregular domain problems. The experimental setup is exactly same as [10]. We experiment with five problems covering both 2D and 3D scenarios, i.e. Irregular Darcy Flow (2D), Pipe Turbulence (2D), Heat Transfer (3D), Composite (3D), and Blood Flow (3D), all of which are defined on irregular physical domains and processed as a set of discrete triangle meshes. The baselines contain GraphSAGE [43], DeepOnet [3], POD-DeepOnet [44], FNO [2] and NORM [10]. We implement the residual learning version of NORM and compare it with the direct learning model under exactly same setup. Appendix B presents more experimental details about the datasets and baselines.
We report the results in Table 1. The residual neural operator NORM (DeltaPhi) consistently outperforms all baselines on all irregular domain problems. The enhancement is particularly significant on Pipe Turbulence and Heat Transfer problem, with 40% - 50% relative improvement (the row "Relative Gain" in Table 1) over original NORM, i.e. the NORM (Previous). These results indicate that the proposed residual operator learning works well on irregular domain problems.
@@
In addition, we visualize the prediction error in Figure 2. The error value is calculated as the absolute difference between the predicted function $\hat{u}$ and ground-truth function $u^{GT}$, i.e. $|\hat{u}-u^{GT}|$. Compared to previous direct learning methods, DeltaPhi significantly reduces the error scale.
@@
Regular domain problems. We conduct the experiment on two regular domain problems i.e. Darcy Flow and Navier-Stokes, following [2]. The training data amount is 100. We validate DeltaPhi on different neural operators, including FNO [2], FFNO [11], CFNO [16], GNOT [23], Galerkin [34], MiOnet [45] and Resnet [42]. The performance is measured with relative error. Appendix B presents more experimental details.
Table 2 presents the experimental results. On both Darcy Flow and Navier-Stokes Equations, the proposed residual operator learning (DeltaPhi) consistently improves performance across different base models. The results validate the effectiveness of DeltaPhi on regular domain problems. However, we observe that the performance improvement of the model on the Navier-Stokes Equation is relatively marginal. We conjecture this is due to the complexity of the fluid dynamic with the viscosity $\upsilon=1e-5$, whereby the correlation between input and output functions with long time intervals is diminished, i.e. similar initial functions $a(x)$ do not necessarily correspond to sufficiently similar output functions $u(x)$. This is a limitation of the current implementation that does not affect the conclusion of this work. It could be alleviated with additional designs such as performing auxiliary trajectory retrieval at each step of time-series prediction to mitigate the decay in correlation between input and output functions.
@@
Model Darcy Flow Navier-Stokes Equation
@@
Table 3: Relative error comparison on different training scales.
Model Different Train Size
@@
Resnet (Previous) 1.25e-1 1.22e-1 1.21e-1 1.21e-1 1.21e-1
Resnet (DeltaPhi) 1.07e-1 9.42e-2 9.08e-2 8.63e-2 8.31e-2
Relative Gain 14.14% $\uparrow$ 22.74% $\uparrow$ 25.14% $\uparrow$ 28.92% $\uparrow$ 31.51% $\uparrow$
Different training scales. We test the residual operator learning on different training set sizes. The experiment is conducted on Darcy Flow problems, with three base models FNO, FFNO, and Resnet. Specifically, we train all models with five training data scales 100, 300, 500, 700, and 900 respectively, and evaluate their performance on the additional same 100 testing samples.
The experimental results are shown in Table 3. On different training set sizes, DeltaPhi improves the prediction performances of neural operators. As the number of available training data decreases, the relative improvement percent (the row "Relative Gain") consistently increases on FNO and FFNO. In addition, several residual neural operators even outperform direct neural operators using less training data amount, e.g. FFNO (DeltaPhi) with training size 500 outperforms FFNO (Previous) with training size 700, Resnet (DeltaPhi) with training size 700 outperforms Resnet (Previous) with training size 900. The experimental results empirically demonstrate the effectiveness of DeltaPhi across different training set scales.
4.2 Residual Learning vs. Direct Learning on Resolution Generalization Problem
@@
Resolution generalization problem. Fourier neural operator [2] enables zero-shot resolution generalization inference. However, when the training data grid is excessively coarse, the inference performance over high resolution data drops a lot. The performance degeneration is more serious when learning operator mapping between fields with weaker spatial continuity such as Darcy Flow.
Setup. We conduct the training-free resolution generalization experiment on Darcy Flow [2]. Specifically, we train the models with low resolutions $85\times 85$, $43\times 43$, $31\times 31$, and $22\times 2$ 22 respectively, then evaluate them using high resolution $421\times 421$ test data. Both FNO [2] and FFNO [11] are compared as the base models. The training set scale includes 100 and 900 trajectories.
@@
Model Train Size=100 Train Size=900
@@
Performance comparison. Table 4 presents the resolution generalization results. Compared to FNO [2], FFNO [11] shows prefer generalization ability for its less number of network weights. In all experimental settings, the proposed physical residual models (tagged with "DeltaPhi") evidently outperform their counterpart base models. The performance improvement of residual operator learning is more noticeable when the training set size is 900. In this setting, FNO (DeltaPhi) managed to match, or even surpass FFNO [11] on resolution $85\times 85$, $43\times 43$, and $22\times 2$ 22, despite using the relatively weaker base network. The results empirically draw the conclusion that physical residual learning improves zero-shot resolution generalization performance.
Training curve comparison. We also report the test loss curve during training in  Figure 3. The training data is 900 and the training resolution is $85\times 85$. As the training proceeds, the relative error on resolution $85\times 85$ data continues to decline. However, after some epochs, the loss on resolution $421\times 421$ gradually increases due to overfitting. The proposed physical residual operator (named with DP-FNO) significantly mitigates this overfitting tendency of the base model (named with FNO).

### v2

@@ 3.3 Auxiliary Sample Retriever $\mathcal{F}^{\text{ret}}$
Inference stage. During inference, given the input function $a^{test}$, we select the most similar auxiliary sample from the training set:
$\mathcal{F}^{\text{ret}}(a^{test},\mathcal{T})=\texttt{argmax}_{j}\,\texttt{Sim}(a^{test},a_{j}).$ (7)
While we take the most similar sample for inference, Section˜A.4 shows that the model is robust to different auxiliary sample choices. When randomly selecting from the top 10 similar samples, repeated testing shows minimal variation - a standard deviation of just 5.41e-5 across five runs. This indicates that neural operators effectively learn to handle varied auxiliary samples for the same input.
Cross-Resolution Sample Retrieval. For cross-resolution scenarios where auxiliary samples have different resolutions from the input function, we employ Fourier Transform-based alignment. Specifically, during inference with high-resolution input function $a^{H}$, we downsample $a^{H}$ to match training resolution via truncating high-frequency components for similarity calculation:
$\mathcal{F}^{\text{ret}}_{\text{aligned}}(a^{H},\mathcal{T})=\mathcal{F}^{\text{ret}}(\text{DownSampling}(a^{H}),a_{j}),$ (8)
where DownSampling is implemented through frequency-domain truncation:
$\text{DownSampling}(a^{H})=\mathcal{F}_{\text{fourier}}^{-1}(\text{Truncate}(\mathcal{F}_{\text{fourier}}(a^{H}))).$ (9)
Here $\mathcal{F}_{\text{fourier}}$ and $\mathcal{F}_{\text{fourier}}^{-1}$ denote the Fourier transform and its inverse, respectively. The Truncate operation removes high-frequency components in Fourier space to match the target resolution. This Fourier-based downsampling preserves the low-frequency structures of the physical field while ensuring consistent resolution for similarity computation. This resolution alignment enables effective utilization of auxiliary samples across different resolutions.
Retrieval Cost Analysis. The computational overhead of our retrieval process is minimal compared to neural network inference. The total complexity consists of similarity score calculation ($O(N_{field}\cdot N_{train})$) and ranking ($O(N_{train}\cdot\log N_{train})$), where $N_{field}$ is the number of field points and $N_{train}$ is the training set size. Our comprehensive experiments on Darcy Flow demonstrate the high efficiency: GPU-based retrieval only takes 0.2-0.3ms across different training set sizes (100-900), and the entire residual learning pipeline, including both retrieval and data preparation, adds merely 0.5ms to the inference time. This negligible overhead makes our method highly practical for real-world applications. Detailed complexity analysis and timing experiments are provided in Section˜C.2.
{algorithm}
[…]
3.4 Architecture of Residual Neural Operators $\mathcal{G}_{\theta}^{\Delta}$
Residual Neural Operator Backbone. Benefiting from the universal approximation capability of neural operator networks (li2020fourier, ; tran2021factorized, ) for direct operator learning, they could be quickly employed as residual neural operators $\mathcal{G}_{\theta}^{\Delta}$ for learning residual operator mapping $\mathcal{G}^{\Delta}$:
$\mathcal{G}_{\theta}^{\Delta}=Q\circ\sigma(W_{l}+\mathcal{K}_{l})\circ\cdots\circ\sigma(W_{1}+\mathcal{K}_{1})\circ P,$ (10)
where $P$ and $Q$ are projection layers, $\mathcal{K}_{i}$ and $W_{i}$ are learnable operators, and $\sigma$ is a nonlinear activation. Next, we introduce how to modify existing neural operators for learning residual mappings.
Concatenation of Auxiliary Samples. As aforementioned, the residual operator $\mathcal{G}^{\Delta}_{\theta}$ requires additional auxiliary function $a_{k_{i}}$ as input. In specific implementation, $a_{i}$ and $a_{k_{i}}$ could be straightly concatenated along their channel dimension:
@@ 3.4 Architecture of Residual Neural Operators $\mathcal{G}_{\theta}^{\Delta}$
where $\oplus$ represents the concatenation operation of two vectors. Then the concatenated vector $a_{input}$ is feed into the neural operator network:
@@
where $P$ is the fully connected encoder in the operator network (Equation˜10), and $v_{0}$ represents the first latent function in the hidden space.
Concatenation of Additional Information. The residual operator mapping can be challenging to learn due to the variability of output residuals for different auxiliary samples. To mitigate such difficulty, we introduce additional information related to $a_{k_{i}}$ as input:
  • Taking the corresponding $u_{k_{i}}(x)$ of $a_{k_{i}}(x)$ as the input.
  • Utilizing calculated relation metrics, e.g. similarity between $a_{i}$ and $a_{k_{i}}$, as inputs.
Our experimental results (Section˜A.3) indicate that such customized inputs could improve the performance of residual neural operators.
Physical Residual Connection. Based on Equation˜4, to obtain the final predicted solution $\hat{u}_{i}$, we introduce a physical residual connection at the final layer of the base operator network. This connection adds the auxiliary solution $u_{k_{i}}$ to the predicted residual:
@@
where $Q$ is the last fully connected projector of the operator network, $v_{l}$ is the last latent function in hidden space produced by $\sigma(W_{l}+\mathcal{K}_{l})$ shown in Equation˜10. This enables end-to-end training with existing frameworks while maintaining consistent setups (loss functions, trainable parameters, etc. ) with the base operator networks.
Cross-Resolution Sample Integration. For cross-resolution scenarios where we need to predict high-resolution outputs $u^{H}$ using low-resolution auxiliary trajectories $(a_{k_{i}},u_{k_{i}})$, we employ Fourier-based upsampling to align the resolutions before integration:
$a_{input}=\text{UpSampling}(a_{k_{i}})\oplus a^{H}.$ (14)
Similarly, for the residual connection, we upsample the low-resolution auxiliary solution:
$\hat{u}^{H}=Q(v_{l})+\text{UpSampling}(u_{k_{i}}),$ (15)
where $Q$ is the decoder and $v_{l}$ is the final hidden state.
The upsampling is performed through zero-padding in Fourier space to preserve the low-frequency structures while extending to higher resolution, formulated as follows:
$\text{UpSampling}(a)=\mathcal{F}_{\text{fourier}}^{-1}(\text{ZeroPad}(\mathcal{F}_{\text{fourier}}(a))),$ (16)
where $\mathcal{F}_{\text{fourier}}$ and $\mathcal{F}_{\text{fourier}}^{-1}$ denote the Fourier transform and its inverse, and ZeroPad extends the spectral representation with zeros to match the target resolution. This Fourier-based approach ensures smooth upsampling while preserving the physical characteristics of the auxiliary trajectories.
3.5 How DeltaPhi Improves Generalization Capability
In data-limited scenarios, DeltaPhi enhances generalization by mitigating overfitting, which is achieved through a form of implicit data augmentation  [^1] We name this implicit data augmentation because there are no more explicit input-output physical function pairs, but directly training on more diverse residual pairs. enabled by our residual learning framework.
Mitigating Overfitting via Implicit Data Augmentation. The core benefit of residual learning is its effectiveness in mitigating overfitting, particularly in data-limited settings. This is achieved through two complementary mechanisms.
First, DeltaPhi functions as an implicit data augmentation strategy. For a training set with $N$ samples, direct learning provides only $N$ input-output pairs. In contrast, DeltaPhi transforms each sample $(a_{i},u_{i})$ into $K$ distinct residual pairs $\{(a_{i},a_{k}),u_{i}-u_{k}\}^{K}_{k=1}$ by pairing it with $K$ different auxiliary samples selected via Section˜3.3. The effectiveness of this augmentation stems from the stability property of PDEs (Section˜3.1), which establishes a structured relationship between input differences and solution residuals. This property allows us to meaningfully control the diversity of the augmented training residuals by sampling auxiliary inputs with varying similarity levels. Therefore, with an opportune retrieval range, we can expand the sparse training distribution—thus mitigating distribution bias—while ensuring the residual learning task remains feasible and does not involve highly dissimilar, difficult-to-learn pairs. This process significantly increases the density and diversity of the training distribution, directly addressing the problem of sparse training data.
Second, the residual learning formulation itself discourages memorization. Because the predicted residual $\mathcal{G}_{\theta}^{\Delta}(a_{i},a_{k_{i}})$ must adapt to different auxiliary samples $a_{k_{i}}$ for the same input $a_{i}$, the model is discouraged from simply memorizing a fixed input-output mapping. Thus, the model is less prone to memorizing labels for specific input functions in data-limited scenarios.
From a geometric perspective, this framework effectively shifts the learning paradigm from approximating isolated points (absolute solutions) on the solution manifold to learning tangent vectors (solution residuals) that describe the manifold’s local geometry. Learning these relational vectors provides a richer learning signal. The stability property discussed in Section˜3.1 ensures this task is tractable, as it establishes a structured relationship between input differences and solution residuals, allowing us to expand the training distribution while maintaining learnability.
Tractability of Residual Mapping Learning. It’s noted that learning the residual mapping $\mathcal{G}^{\Delta}$ can be more difficult than learning the direct mapping $\mathcal{G}$. However, this increased difficulty does not arise from the introduction of spurious high-frequency noise; if two solution functions are smooth, their residual will also be smooth. Instead, the challenge stems from the inherent complexity of the residual mapping itself, which requires the model to synthesize information from two separate input functions ($a_{i}$ and $a_{k_{i}}$). Despite this increased complexity, our experiments demonstrate that existing neural operators are fully capable of learning this mapping effectively, even learning to respect its inherent symmetry without explicit constraints (see Section˜A.8).
Influence of Retrieval Range. While the retrieval range controls the diversity of residuals for training, our results in Section˜A.1 show DeltaPhi is robust to the selection of this hyperparameter. For experimental consistency, we set $K=20$ throughout our comparisons. We further analysis its impact across different test splits in Section˜A.2 and discuss its selection in Section˜C.1. In addition, we introduce an alternative sampling strategy without $K$ selection in Section˜A.6.
@@ 4 Experiment
This section provides the comparison on various PDEs, comparison on resolution generalization problems, and statistical analysis of residual operator learning. We present more experimental results and analysis in Appendix˜A.
4.1 Comparison on Various PDEs
We evaluate residual operator learning across irregular and regular domain PDEs, testing performance with different training data sizes. For fair comparisons, all comparisons use exactly identical hyperparameters, model weights, optimizer settings, and training steps.
Irregular domain problems. Following (chen2023learning, ), we evaluate five problems: 2D (Irregular Darcy Flow, Pipe Turbulence) and 3D (Heat Transfer, Composite, Blood Flow) scenarios, all defined on irregular domains represented by discrete triangle meshes. We compare against GraphSAGE (hamilton2017inductive, ), DeepOnet (lu2019deeponet, ), POD-DeepOnet (lu2022comprehensive, ), FNO (li2020fourier, ) and NORM (chen2023learning, ) baselines, implementing a residual learning version of NORM for direct comparison. Appendix˜B provides more details.
Results in Table˜1 show that NORM (DeltaPhi) outperforms all baselines across irregular domain problems. Performance gains are most notable for Pipe Turbulence and Heat Transfer, showing 40-50% improvement over NORM, demonstrating the residual operator’s effectiveness.
@@ 4.1 Comparison on Various PDEs
In addition, we visualize the prediction error in Figure˜2. The error value is calculated as the absolute difference between the predicted function $\hat{u}$ and ground-truth function $u^{GT}$, i.e. $|\hat{u}-u^{GT}|$. Compared to previous direct learning methods, DeltaPhi significantly reduces the error scale.
Regular domain problems. We conduct the experiment on two regular domain problems i.e. Darcy Flow and Navier-Stokes, following (li2020fourier, ). The training data amount is 100. We validate DeltaPhi on different neural operators, including FNO (li2020fourier, ), FFNO (tran2021factorized, ), CFNO (brandstetter2022clifford, ), GNOT (hao2023gnot, ), Galerkin (cao2021choose, ), MiOnet (jin2022mionet, ) and Resnet (he2016deep, ). The metric is relative L2 error. Appendix˜B presents more details.
Table˜2 shows DeltaPhi improves performance across all base models for both Darcy Flow and Navier-Stokes equations, confirming its effectiveness on regular domains. The improvement on the Navier-Stokes, while modest compared to Darcy Flow, is particularly noteworthy given the inherently chaotic nature of fluid dynamics at viscosity $\upsilon=1e-5$. Even within this challenging scenario characterized by weakened correlation between similar initial conditions $a(x)$ and their evolved states $u(x)$, DeltaPhi still achieves consistent gains. This robustness to complex dynamics strengthens conclusions regarding the method’s general applicability. Performance on temporal systems could be further enhanced through designs such as performing auxiliary sample retrieval at each step of time-series prediction to mitigate the decay in correlation between input and output functions.
@@
Model
Darcy Flow Navier-Stokes Equation
@@
Table 3: Comparison on different training scales.
Model
Different Train Size
@@
Different training scales. We train residual operators using FNO, FFNO, and Resnet on Darcy Flow with varying training set sizes (100, 300, 500, 700, and 900) and test their performance on the additional 100 samples.
The experimental results are shown in Table˜3. On different training set sizes, DeltaPhi improves the prediction performances of neural operators. As the number of available training data decreases, the relative improvement percent (the row "Relative Gain") consistently increases on FNO and FFNO. In addition, several residual neural operators even outperform direct neural operators using less training data amount, e.g. FFNO (DeltaPhi) with training size 500 outperforms FFNO (Direct) with training size 700, Resnet (DeltaPhi) with training size 700 outperforms Resnet (Direct) with training size 900. The results empirically demonstrate the effectiveness of DeltaPhi across different training set scales.
4.2 Comparison on Resolution Generalization Problem
@@
Resolution generalization problem. Fourier neural operator (li2020fourier, ) enables zero-shot resolution generalization inference. However, when the training data grid is excessively coarse, the inference performance over high resolution data drops a lot. The performance degeneration is more serious when learning operator mapping between fields with weaker spatial continuity such as Darcy Flow.
Setup. We conduct the training-free resolution generalization experiment on Darcy Flow (li2020fourier, ). Specifically, we train the neural operators with low resolutions $85\times 85$, $43\times 43$, $31\times 31$, and $22\times 22$ respectively, then evaluate them using high resolution $421\times 421$ test data. Both FNO (li2020fourier, ) and FFNO (tran2021factorized, ) are compared as the base models. The training set scale includes 100 and 900 trajectories.
@@
Model
Train Size=100 Train Size=900
@@
Performance comparison. Table˜4 presents the resolution generalization results. Compared to FNO (li2020fourier, ), FFNO (tran2021factorized, ) shows preferable generalization ability for its less number of network weights. In all experimental settings, the proposed physical residual models (tagged with "DeltaPhi") evidently outperform their counterpart base models. The performance improvement of residual operator learning is more noticeable when the training set size is 900. In this setting, FNO (DeltaPhi) managed to match, or even surpass FFNO (tran2021factorized, ) on resolution $85\times 85$, $43\times 43$, and $22\times 22$, despite using the relatively weaker base network. The results empirically draw the conclusion that physical residual learning improves zero-shot resolution generalization performance.
Training curve comparison. We also report the test loss curve during training in  Figure˜3. The training amount is 900 and the training resolution is $85\times 85$. As the training proceeds, the relative error on resolution $85\times 85$ data continues to decline. However, after some epochs, the loss on resolution $421\times 421$ gradually increases due to overfitting. The proposed residual operator (named with DP-FNO) significantly mitigates this overfitting tendency of the base model (named with FNO).